40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 9 01:29:26 2017, gautam, Update, LSC, DRMI Nosie Budget v3.0 7x
    Reply  Mon Oct 9 11:55:01 2017, Koji, Update, LSC, DRMI Nosie Budget v3.0 
       Reply  Mon Oct 9 22:18:34 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_dark.pdf
          Reply  Tue Oct 10 22:04:06 2017, rana, Update, LSC, AS55Q Dark Noise 
          Reply  Wed Oct 11 14:42:03 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
             Reply  Wed Oct 11 19:31:32 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
                Reply  Thu Oct 12 01:50:11 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_dark.png
                   Reply  Thu Oct 12 12:17:28 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_AS55_40m.pdf
                      Reply  Fri Oct 13 12:26:12 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_DemodNoises.pdf
                         Reply  Mon Oct 16 16:01:04 2017, gautam, Update, LSC, AS55Q Dark Noise DemodBoardwOscope.pdfDemodBoardwOscope_ASD.pdf
                            Reply  Tue Oct 17 19:31:53 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_TFs.pdfAS55_DemodNoises.pdf
    Reply  Tue Nov 7 17:45:05 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 SRCL_MICH_whitenoise_tf.pdfPRCL_MICH_whitenoise_tf.pdfMICH_aux.pdfC1NB_disp_40m_MICH_NB_2017-10-08.pdf
       Reply  Tue Nov 7 22:56:21 2017, gautam, Update, LSC, DRMI locking recovered DRMI_7Nov20178.png
          Reply  Wed Nov 8 00:28:16 2017, gautam, Update, LSC, Laser intensity coupling measurement attempt PSL_RIN.pdf
       Reply  Wed Nov 8 09:37:45 2017, rana, Update, LSC, DRMI Nosie Budget v3.1 
          Reply  Wed Nov 8 09:59:12 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 
             Reply  Wed Nov 15 01:37:07 2017, gautam, Update, LSC, DRMI low freq. nosie improved C1NB_disp_40m_MICH_NB_2017-11-15.pdfOL_A2L_couplings.pdf
                Reply  Thu Nov 16 00:53:26 2017, gautam, Update, LSC, DRMI noise sub-budgets C1NB_a2l_40m_MICH_NB_2017-11-15.pdfC1NB_aux_40m_MICH_NB_2017-11-15.pdf
Message ID: 13414     Entry time: Wed Nov 8 00:28:16 2017     In reply to: 13413
Author: gautam 
Type: Update 
Category: LSC 
Subject: Laser intensity coupling measurement attempt 

I tried measuring the coupling of PSL intensity noise by driving some broadband noise bandpassed between 80-300Hz using the spare DAC channel at 1Y3 that I had set up for this purpose a couple of weeks ago (via a battery powered SR560 buffer set to low-noise operation mode because I'm not sure if the DAC output can drive a ~20m long cable). I was monitoring the MC2 TRANS QPD Sum channel spectrum while driving this broadband noise - the "nominal" spectrum isn't very clean, there are a bunch of notches from a 60Hz comb and a forest of peaks over a broad hump from 300Hz-1kHz (see Attachment #1).

I was able to increase the drive to the AOM till the RIN in the band being driven increased by ~10x, and saw no change in the MICH error signal spectrum [see Attachment #1] - during this measurement, the RFPD whitening was turned on for REFL11, REFL55 and AS55, and the ITM coil drivers were de-whitened, so as to get a MICH spectrum that is about as "low-noise" as I've gotten it so far.

I tried increasing the drive further, but at this point, started seeing frequent MC locklosses - I'm not convinced this is entirely correlated to my AOM activities, so I will try some more, but at the very least, this places an upper bound on the coupling from intensity noise to MICH.

Attachment 1: PSL_RIN.pdf  69 kB  | Show | Show all
ELOG V3.1.3-